a common stock will have a price of either $75 or $25 in two months. a two month call option on the stock has a strike price of $65. create a riskless portdolio. sell 100 call options, how many shares should you buy?
Answer #1

Strike price =

65

Upside value =

75

Downside value =

25

VC at upper = Upper value - Strike price (subject to 0)

75 - 65 = 10

VC at down

25 - 65 = 0 (it cannot be negative)

Hedge ratio = (V.C. at Upper value -V.C. at Down value) / (Upper Value - Down value)

(10-0) /(75 - 25)

0.2

Hedge ratio is 0.2. It means if one share is long then 0.20 call should be short.

Formula for no. of shares buy = no. of call options/Hedge ratio

so no.of shares buy should be = 100/0.20

500

So, 500 shares should be bought to create riskless portfolio.